# On the Noisy Road to Open Quantum Dynamics: The Place of Stochastic Hamiltonians

### v1

@@ III Unitary propagation under stochastic Hamiltonians
Looking for the solution of eq. 6, we expect to be able to write a formal propagator of the form
@@
As anticipated in the introduction, we want to write eq. 6 in a formally correct way. We start writing the random potential as a noise, for simplicity the common white noise $\xi_{t}$ of intensity $\sigma$, by setting
$Z_{t}=\xi_{t}=\sigma\frac{\mathop{}\!\mathrm{d}{W_{t}}}{\mathop{}\!\mathrm{d}{t}},$ (19)
@@
$\mathop{}\!\mathrm{d}\psi_{t}=-i{H}\psi_{t}\mathop{}\!\mathrm{d}{t}-i\sigma{R}\psi_{t}\circ\mathop{}\!\mathrm{d}{W_{t}}.$ (20)
The choice of the Stratonovich interpretation of the differential is fundamental, as it allows us to write the analytical solution to eq. 20, since the rules of ordinary calculus are valid, in the same fashion as for the solution of the Schrödinger equation
$\psi_{t}=\mathcal{T}_{+}\exp\left\{-i\int_{0}^{t}H\mathop{}\!\mathrm{d}{t}-i\sigma\int_{0}^{t}R\circ\mathop{}\!\mathrm{d}{W_{t}}\right\}\psi_{0}\,$, (21)
where we recognize in the time-ordered exponential the propagator of the dynamics, meaning that each trajectory can be propagated through this unitary. These solutions, proposed using white noise for the sake of clarity, can be easily generalized for colored noise, paying attention to the correct integration of the deterministic and stochastic parts.
Care must be used in the description of the dynamics of the mean density matrix of the open system, whose equation cannot be obtained by sticking to this formalism. Indeed, while eq. 7 holds true, if we take the product differential $\mathop{}\!\mathrm{d}{(\psi_{t}\psi_{t}^{\dagger})}$ and we average,
$\mathbb{E}\left[\mathop{}\!\mathrm{d}{(\psi_{t}\psi_{t}^{\dagger})}\right]=-i\left[H,\psi_{t}\psi_{t}^{\dagger}\right]\mathop{}\!\mathrm{d}{t}-i\mathbb{E}\left(\left[R,\psi_{t}\psi_{t}^{\dagger}\right]\circ\mathop{}\!\mathrm{d}{W_{t}}\right)$ (22)
@@ III.1 Use of unitary propagators for implementation of quantum algorithms
In the framework discussed above, we obtain the expression of the propagator as a unitary operator, and norm-preservation is thus ensured trajectory-wise. We can obtain the numerical solution for the evolution in the time interval $[0,T]$, discretizing time into $S$ time steps $\tau=T/S$ and evolving the initial state through the first-order Magnus expansion [Magnus1954OnOperator, Blanes2009TheApplications] of the time-ordered propagator
$\mathcal{U}_{t}\approx\ \prod_{s=1}^{S}\ \mathcal{U}_{s\tau,(s-1)\tau}$ (23)
where each time-step unitary propagator is
@@
$\begin{split}\mathcal{U}_{s\tau,(s-1)\tau}=\exp\Bigg\{&-i\int_{(s-1)\tau}^{s\tau}H\mathop{}\!\mathrm{d}{r}\\ &-i\sigma\int_{(s-1)\tau}^{s\tau}R\circ\mathop{}\!\mathrm{d}{W_{r}}\Bigg\}\end{split}$ (24)
@@
and allows for numerical integration of every trajectory solution. For a time-independent system Hamiltonian and noise operator, the short-time evolution operators read
$\mathcal{U}_{\tau}=\exp\left\{-iH\tau-i\sigma R\Delta{W_{\tau}}\right\},$ (25)
@@
where the white noise increment $\Delta{W_{\tau}}$ is drawn from a normal distribution of zero mean and standard deviation the square root of the time interval:
$\Delta{W\tau}\sim\mathcal{N}(0,\sqrt{\tau}).$ (26)
This numerical scheme, also used for implementations on classical computers [Jansen2006, Liang2012], allows the construction of quantum algorithms in digital quantum computer architectures without the overhead of auxiliary qubits. A simple example of a quantum algorithm with classical noise [gallina2022, Gallina2024FromDynamics] is depicted in LABEL:fig:circuit_QCNA, for a qubit evolution with Hamiltonian $H=\varepsilon\sigma_{z}+\Omega\sigma_{x}$, implemented using its first-order Suzuki-Trotter approximation [Trotter1959OnOperators, Suzuki1976GeneralizedProblems], and noise $\xi_{t}$ acting on the couplings through $R=\sigma_{x}$ (where $\sigma_{x}$ and $\sigma_{z}$ are the standard Pauli matrices).

### v2

@@ III Unitary propagation under stochastic Hamiltonians
Looking for the solution of eq. 6, under ordinary calculus rules, we expect to be able to write a formal propagator of the form
@@
As anticipated in the introduction, we want to write eq. 6 in a formally correct way. We start writing the random potential as a noise, for simplicity the common white noise $\xi_{t}$ of intensity $\gamma$, by setting
$Z_{t}=\xi_{t}=\gamma\frac{\mathop{}\!\mathrm{d}{W_{t}}}{\mathop{}\!\mathrm{d}{t}},$ (21)
@@
$\mathop{}\!\mathrm{d}\psi_{t}=-i{H}\psi_{t}\mathop{}\!\mathrm{d}{t}-i\gamma{R}\psi_{t}\circ\mathop{}\!\mathrm{d}{W_{t}}.$ (22)
The choice of the Stratonovich interpretation of the differential is fundamental, as it allows us to write the analytical solution to eq. 22, since in this formalism the rules of ordinary calculus are indeed valid. As for the solution of the Schrödinger equation, we obtain
$\psi_{t}=\mathcal{T}_{+}\exp\left\{-i\int_{0}^{t}H\mathop{}\!\mathrm{d}{t}-i\gamma\int_{0}^{t}R\circ\mathop{}\!\mathrm{d}{W_{t}}\right\}\psi_{0}\,$, (23)
where we recognize in the time-ordered exponential the propagator of the dynamics, meaning that each trajectory can be propagated through this unitary. These solutions, proposed using white noise for the sake of clarity, can be easily generalized for colored noise, paying attention to the correct integration of the deterministic and stochastic parts. We point out, in the following LABEL:sec:ito, that this is not always the case in the Itō formalism, and further constraints must be introduced.
Care must be used in the description of the dynamics of the mean density matrix of the open system, whose equation cannot be obtained by adhering to this formalism. Indeed, while eq. 7 holds true, if we take the product differential $\mathop{}\!\mathrm{d}{(\psi_{t}\psi_{t}^{\dagger})}$ and we average,
$\mathbb{E}\left[\mathop{}\!\mathrm{d}{(\psi_{t}\psi_{t}^{\dagger})}\right]=-i\left[H,\psi_{t}\psi_{t}^{\dagger}\right]\mathop{}\!\mathrm{d}{t}-i\gamma\,\mathbb{E}\left(\left[R,\psi_{t}\psi_{t}^{\dagger}\right]\circ\mathop{}\!\mathrm{d}{W_{t}}\right)$ (24)
@@ III.1 Use of unitary propagators for implementation of quantum algorithms
In the framework discussed above, we obtain the expression of the propagator as a unitary operator, and norm-preservation is thus ensured trajectory-wise. We can obtain the numerical solution for the evolution in the time interval $[0,T]$, discretizing time into $S$ small time steps $\tau=T/S$
$\mathcal{U}_{t}=\ \prod_{s=1}^{S}\ \mathcal{U}_{s\tau,(s-1)\tau}$ (25)
and approximating at each time step the time-ordered unitary propagator with a first-order Magnus expansion [Magnus1954OnOperator, Blanes2009TheApplications]:
@@
$\begin{split}\mathcal{U}_{s\tau,(s-1)\tau}\approx\exp\Bigg\{&-i\int_{(s-1)\tau}^{s\tau}H\mathop{}\!\mathrm{d}{r}\\ &-i\gamma\int_{(s-1)\tau}^{s\tau}R\circ\mathop{}\!\mathrm{d}{W_{r}}\Bigg\}.\end{split}$ (26)
This allows for the numerical integration of every trajectory solution, keeping the unitary character of the propagators. We highlight that the choice of a small time step is important for convergence to the exact time-ordering solution. Higher-order Magnus expansions can also be employed to achieve better convergence.
@@
For a time-independent system Hamiltonian and noise operator, the short-time evolution operators read
$\mathcal{U}_{\tau}=\exp\left\{-iH\tau-i\gamma R\Delta{W_{\tau}}\right\},$ (27)
@@
where the white-noise increment $\Delta{W_{\tau}}$ is drawn from a normal distribution of zero mean and standard deviation the square root of the time interval:
$\Delta{W\tau}\sim\mathcal{N}(0,\tau)=\sqrt{\tau}\,\mathcal{N}(0,1).$ (28)
This numerical scheme, also used for implementations on classical computers [Jansen2006, Liang2012], allows the construction of quantum algorithms for digital quantum computers without the overhead of auxiliary resources (auxiliary qubits). This is because the unitary character ensured by the use of the Magnus expansion is naturally suited to be implemented in the language of quantum gates, the unitary operations that represent the building blocks of every quantum circuit [Nielsen2010QuantumEdition].
Example 2 (Quantum Circuit) A simple example of a quantum algorithm [gallina2022, Gallina2024FromDynamics] that implements eqs. (25) and (27) is depicted in LABEL:fig:circuit_QCNA. The quantum circuit reproduces the evolution of a two-level system (qubit) with states $|0\rangle$ and $|1\rangle$) dictated by a system Hamiltonian $H=\varepsilon\sigma_{z}+\Omega\sigma_{x}$, where $\sigma_{z}=|0\rangle\langle 0|-|1\rangle\langle 1|$ and $\sigma_{x}=|0\rangle\langle 1|+|1\rangle\langle 0|$, and a fluctuation Hamiltonian $\hat{H}_{t}^{\mathrm{fluct}}=\xi_{t}\sigma_{z}$. This is formally equivalent to the model used to represent the coupled excited states in section II.1, cast in terms of Pauli operators acting on a qubit. Explicitly, the energy difference of the sites is expressed by $\varepsilon=(\varepsilon_{0}-\varepsilon_{1})/2$ and so the fluctuation $\xi_{t}=(\delta\varepsilon_{0}(t)-\delta\varepsilon_{1}(t))/2$. The evolution operators, described in eq. (27), can be implemented in terms of quantum gates (involving 1 or 2 qubits) in various ways. Here, for illustrative purposes, we show the notorious first-order Suzuki-Trotter decomposition [Trotter1959OnOperators, Suzuki1976GeneralizedProblems], which allows us to write
$\begin{split}&\mathcal{U}_{\tau}=\exp\{-i\varepsilon\tau\sigma_{z}-i\Omega\tau\sigma_{x}-i\gamma\Delta W_{\tau}\sigma_{z}\}\\ &\approx\left[\exp\left\{-i\frac{\gamma\Delta W_{\tau}}{n}\sigma_{z}\right\}\exp\left\{-i\frac{\Omega\tau}{n}\sigma_{x}\right\}\exp\left\{-i\frac{\varepsilon\tau}{n}\sigma_{z}\right\}\right]^{n}\end{split}$ (29)
where, in general, the larger $n$ the better the convergence. However, due to the choice of a small time step dictated by the Magnus expansion, good results can be obtained even with a small $n$ (e.g., $n=1$). Each unitary operator inside the squared brackets in eq. (29) is represented by a gate $R_{\alpha}(2\theta)=\exp(-i\theta\sigma_{\alpha})$, with $\alpha\in\{Z,X\}$, inside the red dashed barriers in LABEL:fig:circuit_QCNA. To compute the dynamics of the system up to time $T=S\tau$, the 3-gate sequence is repeated $nS$ times. At the end, the qubit ($q_{0}$) is measured to know the population of the two states, $|0\rangle$ and $|1\rangle$.